# Atlas: A Framework for ML Lifecycle Provenance & Transparency

### v1

@@ 3.3.1 Trusted Parties
As mentioned, Atlas considers the transparency services, verifiers and model users in an ML lifecycle to be trusted. We make the following assumptions about the security of the systems underlying and enabling the participants: 1) the hardware and cryptographic primitives are implemented correctly and do not contain known vulnerabilities; 2) a separate PKI system and organizations representing the participants follow best practices for key management, network security and access control.
This allows us to trust that transparency services correctly implement Atlas attestation mechanisms, detect tampering with them by a malicious MLaaS provider and maintain accurate golden values of model artifacts. Verifiers are trusted to properly evaluate Atlas attestations, including their digital signatures, against pre-specified expectations, and to publish a record of their evaluations.
@@ 3.3.2 Out of Scope
We consider the following threats to be out of scope for Atlas. While a critical threat to deployed AI applications, inference time black-box attacks (e.g., evasion attacks, model extraction, membership inference) are caused by malicious users, which we do not address; solutions proposed to reduce this particular risk [45, 46, 47] are complementary to our work.
Side-channel attacks against hardware enclaves, physical attacks on hardware infrastructure, and network-level denial of service attacks are also beyond the scope of Atlas. These attacks are the subject of a large body of prior work [36], and we acknowledge that these complementary security measures could be added to deployments of Atlas.
@@ 3.4 Design Requirements
R1: Artifact tampering is detectable. To provide model artifact authenticity, Atlas must enable verifiers to detect unexpected modifications to model artifacts by malicious MLaaS providers or hubs.
R2: Every model transformation is attested. Because adversaries may seek to tamper with model artifacts after they are produced by a pipeline, every model transformation in Atlas must produce an authenticated record as evidence for the process.
R3: Verifiable model lineage. Atlas verifiers must be able to detect unauthorized operators and unintended or malicious changes to the expected stages of the lifecycle (e.g., pipelines operating out of order, or being omitted).
R4: Strongly isolated ML systems. To reduce the risks of tampering with a pipeline during its execution, Atlas must restrict access to its ML system by unauthorized entities, and contain compromises from propagating beyond the execution environment.
R5: Controlled artifact disclosure. Atlas must address data and model producer concerns about PII and intellectual property confidentiality by enabling participants to control which parties and systems are able to access datasets and model components.
R6: Pipeline agnostic. To facilitate adoption, Atlas must be agnostic to any ML pipeline that integrates it.
R7: Efficiency. We seek to minimize the computational and storage overheads incurred by Atlas to enable the implementation and deployment of Atlas in ML systems using commodity platforms and services.
@@ 4 Atlas Framework
Atlas introduces two core components to the ML lifecycle: 1) the transparency service interacting with MLaaS providers; 2) the verification service that model users and verifiers use to verify integrity and provenance. Fig. 1 depicts an example ML model lifecycle with Atlas.
Figure 1: Example ML lifecycle with Atlas. Striped boxes and bolded arrows represent Atlas components and data flows. MLaaS providers obtain model artifacts (datasets or models) and perform transformations on them. Atlas attestation clients generate and submit digitally signed artifact and transformation metadata to a transparency log. Model users commonly delegate model provenance checking to an Atlas verifier.
4.1 Transparency Service
To provide transparency across stages of the ML lifecycle, MLaaS providers integrate an Atlas attestation client in their system. The transparency log makes all attestation clients’ collected metadata available to verifiers. The four core techniques underlying the Atlas transparency service are designed to be general, allowing them to remain agnostic to the particular ML lifecycle stage or pipeline they are applied to (R6).
@@ 4.1.1 Artifact Measurements
The Atlas attestation client within an MLaaS provider cryptographically measures every artifact ingested into and output by an ML pipeline. That is, we use a collision resistant cryptographic hash function to compute unique, immutable measurement values for every artifact. Because any changes to a given artifact result in a different hash value, Atlas verifiers wishing to check whether a given model artifact matches their expectation are able to detect tampering between stages of the lifecyle (R1).
@@ 4.1.2 Model Transformation Integrity
In addition to capturing measurements about all inputs and outputs of an ML pipeline, the Atlas attestation client continuously verifies new inputs and monitors an ML system’s execution to collect detailed information about pipeline state and operations that transform the input artifacts. That is, during data processing, the client tracks dataset modifications and preprocessing operations; during model training, Atlas captures state changes in model weights, hyperparameters, and configurations.
To enhance the integrity of the ML artifacts and system at runtime, Atlas runs ML pipeline code within trusted execution environments (TEEs) (e.g., Intel TDX [48] or AMD SEV-SNP [49]), which serve as a hardware-based root of trust for ML system component measurements extending from its hardware to the firmware and software.
These measures allow the Atlas attestation client to verify the integrity and authenticity of the compute environment before beginning the execution of the pipeline. Throughout execution, the TEE maintains isolated memory regions, reducing the risk of interference with the pipeline, as well as unauthorized disclosure of artifacts and pipeline code, by a compromised MLaaS provider (R4 & R5).
At the conclusion of an operation in a given pipeline, the client generates a record with all collected ML system information, and digitally signs it with a TEE-generated key along with all artifact measurements, providing a model transformation attestation (R2). If required by the artifact producer, Atlas may encrypt the produced model artifacts prior to digitally signing and uploading them to a given hub to further enhance artifact authenticity and confidentiality while in transit and at rest (R1 & R5).
@@ 4.1.3 Provenance Chains
To allow verifiers to trace an ML model throughout all of its lifecycle stages from data preparation to deployment, Atlas embeds the cryptographic hash of the transformation attestation for every pipeline input that passed verification into the output attestation. These hash values are digitally signed as part of the transformation attestation. Thus, Atlas attestation clients establish an authenticated, cryptographically verifiable provenance chain representing a model’s lineage relationships throughout its lifecycle (R3).
4.1.4 Transparency Log
@@
On the server side, a transparency log contains the golden values (i.e., known good values) of a producer’s model artifacts, an MLaaS provider’s system components, and the transformation attestations collected by the clients throughout the ML lifecycle. To enable efficient insertion and provenance verification while accommodating the cyclical nature of the ML lifecycle, Atlas relies on two data structures.
@@
First, to provide cryptographic tamper-evidence for the stored values, the transparency log is constructed using an append-only Merkle tree [50], meaning that pipeline metadata can be efficiently inserted in the right-most empty leaf node of the tree (e.g., as in [28]). Second, to enable more efficient verification of provenance across pipelines (or even cycles of the ML lifecycle), Atlas can represent each discreet pipeline/cycle using a different Merkle tree and chain the trees by embedding the Merkle root hash of the upstream pipeline or previous cycle into the “current” Merkle tree’s root (e.g., as in [51]), limiting tree traversal operations when verifying a lifecycle over time (R7).
We note that the choice of whether and how to partition the transparency log’s Merkle trees will result in different security and performance tradeoffs; depending on the party that operates the transparency service in practice, they may favor maintaining a smaller or larger number of trees according to their specific metadata access control and resource requirements.
4.2 Verification System
The Atlas verification system performs staged verification. When a pipeline ingests an input artifact, the attestation client requests the artifact’s verification at the verification service, interacting with the transparency log to obtain the golden values and provenance information.
First, the verification service validates that the received metadata was produced by the expected parties via their digital signatures, and checks whether the artifact matches the expected golden value measurements. If the first verification stage passes, Atlas inspects the received transformation metadata to confirm the recorded ML system and pipeline operations ran as expected based on their TEE-backed attestations. Third, the artifact’s lineage is validated by tracing its provenance chain to check its transformations as far back as specified in the verifier’s policy.
To enable efficient batch verification of input artifacts, Atlas groups related artifacts including training datasets, model weights, algorithm implementations, and evaluation results. By maintaining a cache of verified states and an incremental verification flow, the system avoids re-computing cryptographic operations for unchanged artifacts (R7). This particularly benefits ML pipelines where only small portions of the workflow change between iterations.
@@ 5 Implementation
Our proof-of-concept implementation integrates with Kubeflow [33] through standard APIs for metadata tracking and execution monitoring within ML pipelines. This integration approach enabled us to avoid significant modifications to our case study pipelines (§6), while maintaining Atlas’ security and transparency enhancements.
The Atlas attestation client is implemented as two components. First, the metadata sidecar (§5) captures ML artifact and pipeline metadata in C2PA manifest format [10]. Due to current poor support for automated C2PA manifest generation and verification for ML models, we implemented Rust-based library in 2.3k LoC that captures cryptographic measurements, TDX attestation hashes, digital signatures, and temporal metadata. Second, the attestation client implementation leverages a dedicated TEE service (§5, which interacts with Intel TDX [48], a virtualization-layer TEE, to provide the hardware-based security primitives for Atlas.
To implement the transparency log, our framework extends the Sigstore’s Rekor  [28] tamper-evident ledger. Building on Rekor’s existing capabilities, we added support for storage and verification of Atlas C2PA-based model transformation attestations. The modified implementation validates C2PA signatures and attestation measurements during manifest submission, ensuring only properly signed and attested artifacts are stored in the log.
For ease of implementation, our prototype verification system (§5.3) is integrated into the client’s metadata sidecar.
Metadata Sidecar. To integrate our C2PA tooling non-intrusively with ML pipelines, the Atlas attestation client uses a sidecar container architecture. This allows the client to monitor the pipeline for modifications, generate manifests and update relationships through an optimized storage layer. In addition to collecting C2PA manifest metadata, the sidecar generates the authenticated transformation attestations for model artifacts, and tracks artifact relationships by mapping datasets to checkpoints, checkpoints to models, and maintaining configuration bindings.
The sidecar container implementation monitors the pipeline for modifications, generating manifests that are first stored in a local cache layer before being committed to the transparency log. The local cache maintains an indexed hierarchy of manifests for efficient validation during pipeline execution, before the final storage in Rekor for tamper-evident provenance tracking. For storage optimization, our implementation decomposes manifests into constituent components. The assertion store, claim signatures, and metadata are stored separately, with relationships maintained through a reference system. This approach enables efficient updates to specific manifest components, reduced storage redundancy, optimized query performance, and scalable version tracking.
@@
TEE Service. To obtain the Intel TDX-based compute environment attestations that are included in the C2PA metadata, the sidecar interfaces with the TEE service. We implement a dedicated RESTful service that captures TEE state and validates runtime ML system component measurements, which are cryptographically anchored in hardware via Intel TDX. The TEE service implements the Confidential Containers (CoCo) Attestation format [52] for standardized claims used in TEE verification.
We optimize attestation operations through batching and caching mechanisms.The client coordinates with the verification system through standard Kubeflow interfaces, enabling seamless integration with the other Atlas components while preserving attestation integrity.
5.1 Kubeflow Integration

### v2

@@ 3.3.1 Trusted Parties
Atlas considers the model users, transparency and verification services in an ML lifecycle to be trusted. We make the following assumptions about the systems supporting these stakeholders: 1) the hardware and cryptographic primitives are implemented correctly and do not contain known vulnerabilities; 2) a separate PKI system exists and organizations representing the stakeholders follow best practices for key management, network security and access control.
Further, running attestation clients in TEEs allows us to trust Atlas metadata generation, or detect attempts of tampering by malicious MLaaS providers. Similarly, model users and verification services can trust the integrity of golden values and attestations stored in Atlas transparency logs (or detect tampering) via their tamper-evident construction (see §4.2). Verification services are trusted to properly evaluate attestations, including their digital signatures, against pre-specified model user expectations.
@@ 3.3.2 Out of Scope
Many available TEEs provide confidentiality features, but addressing PII and model intellectual property concerns end-to-end requires a more comprehensive framework (e.g.,  [61, 7]). We plan to explore confidentiality within Atlas as future work.
While a critical threat to deployed AI applications, Atlas does not address inference time black-box attacks (e.g., evasion attacks, model extraction, membership inference) caused by malicious users; solutions to reduce this risk [62, 63, 64] are complementary.
Side-channel attacks against hardware enclaves, physical attacks on hardware infrastructure, and network-level denial of service attacks are also beyond the scope of Atlas. These attacks are the subject of a large body of prior work [47], and these complementary security measures could be added to deployments of Atlas.
@@ 3.4 Design Requirements
R1: Artifact tampering is detectable. To provide model artifact integrity, Atlas must enable verification services to detect unexpected modifications to model artifacts.
R2: Every model transformation is attested. Because adversaries may seek to tamper with model artifacts after they are produced by a pipeline, Atlas attestation clients must record every model transformation in authenticated metadata as evidence for the process, including all inputs to the transformation.
R3: Verifiable model lineage. Atlas verification services must be able to detect unintended/malicious changes by MLaaS providers to the expected stages of the lifecycle (e.g., pipelines operating out of order, or being omitted), from initial data processing through model deployment.
R4: Strongly isolated ML systems. To detect tampering with a pipeline during its execution, Atlas must restrict access to its ML system by malicious MLaaS providers, and contain compromises from propagating beyond the execution environment.
R5: Pipeline agnostic. To facilitate adoption, Atlas must be agnostic to any ML pipeline that integrates it.
R6: Efficiency. We seek to minimize the computational and storage overheads incurred by Atlas to enable the implementation and deployment of Atlas in ML systems using commodity platforms and services.
@@ 4 Atlas Framework
Atlas introduces two core components to the ML lifecycle: 1) the transparency service interacting with MLaaS providers; 2) the verification service for validating model integrity and provenance.
The core techniques underlying the transparency and verification services are designed to be general, allowing them to remain agnostic to the particular ML lifecycle stage or pipeline they are applied to (R5). Fig. 1 depicts an example ML model lifecycle with Atlas.
Figure 1: Atlas workflow for ML lifecycle transparency in a BERT Meta [11, 12] fine-tuning process. The attestation client monitors ML systems running in TEEs throughout the pipeline stages, collecting provenance metadata from initial deployment through verification.
4.1 Attestation Client
Atlas combines mechanisms for artifact and runtime environment integrity to provide transparency across the stages of the ML lifecycle. Thus, MLaaS providers integrate an Atlas attestation client with an ML system, each running within a dedicated trusted execution environment (TEE) (e.g., Intel TDX [10] or AMD SEV-SNP [65]).
TEEs serve two purposes in Atlas. First, TEE hardware-enforced memory integrity detects runtime tampering with the attestation client and ML system components by the MLaaS host (including privileged software like the OS or hypervisor). Second, TEEs provide a hardware-based root of trust for provenance and pipeline metadata.
@@ 4.1.1 Artifact Measurements
For every artifact that is ingested into and output by an ML pipeline, the attestation client computes a cryptographic measurement using a collision resistant hash function resulting in a unique, immutable identifier. If an artifact is tampered with, its measurement will differ from its golden value, allowing Atlas verification services to detect modifications between lifecycle stages (R1). Artifact producers are expected to publish digitally signed measurements as golden values whenever an artifact is first created.
@@ 4.1.2 Model Transformation Integrity
The attestation client is also responsible for generating provenance and pipeline metadata describing the transformation process and ML system that produced a new model artifact. Before pipeline execution begins, the client obtains a hardware attestation to its initial state from its TEE, which includes measurements of the client’s execution environment.
At the start of pipeline execution, the attestation client remotely attests the ML system’s TEE [10, 65], verifying the integrity of the pipeline’s compute environment. Specifically, the client checks that the ML system’s firmware, OS and pipeline code match the golden values published by the MLaaS provider to ensure that a pipeline starts from a known good state. [^3] Because only select TEE implementations [66] provide built-in support for attested interactions with I/O devices or ML accelerators like GPUs, it is challenging to distinguish between benign and malicious runtime modifications to pipelines via the network, disk, etc (see §6.4).
Throughout pipeline execution, TEE hardware enforces memory integrity checks and isolation of executing attestation client/ML system code and in-memory data, reducing the risks of interference by compromised MLaaS providers or any co-located ML systems (R4). Further, the attestation client continuously monitors an ML system’s execution, which allows it to determine when artifacts move into or outside of the system, and to collect information about operations that transform the input artifacts (see §5). For example, during data processing, the client tracks dataset modifications and preprocessing operations; during model training, Atlas captures state changes in model weights, hyperparameters, and configurations.
When pipeline execution concludes, the attestation client generates pipeline metadata containing all collected ML system runtime information and the ML system’s TEE attestation. Then, the client creates provenance metadata including 1) artifact measurements, 2) operations producing the outputs, 3) TEE attestation for the client, 4) pipeline metadata.
The attestation client then digitally signs the provenance metadata with keys it generates within the TEE, cryptographically binding ML artifacts to the pipeline and precursor artifacts that created them in a transformation attestation (R2). The client uploads this attestation to the transparency service.
@@ 4.1.3 Provenance Chains
To enable ML model provenance tracking throughout all of its lifecycle stages, the attestation client embeds the cryptographic hash of precursor artifact attestations into every artifact’s transformation attestation. These hash values are digitally signed as part of the transformation attestation, enabling detection of unexpected/malicious modifications between attested ML artifact transformations. Thus, Atlas attestation clients establish an authenticated, verifiable provenance chain representing a model’s lineage relationships (R3).
4.2 Transparency Log
The transparency service’s log makes all published golden values and attestation client-generated metadata available to verification services.
@@ 4.2 Transparency Log
To enable efficient insertion and provenance verification while accommodating the cyclical nature of the ML lifecycle, Atlas relies on two data structures (R6).
@@
First, to provide cryptographic tamper-evidence for the stored values, the transparency log is constructed using an append-only Merkle tree [9], meaning that pipeline metadata can be efficiently inserted in the right-most empty leaf node of the tree (e.g., as in [32]). Second, to enable more efficient verification of provenance across pipelines (or even cycles of the ML lifecycle), Atlas can represent each discrete pipeline/cycle using a different Merkle tree. These separate trees are linked by embedding the Merkle root hash of the preceding pipeline or cycle into the “latest” Merkle tree structure (e.g., as in [67, 68]), providing a temporal cryptographic tree chain. Optimizations to chained Merkle trees have been developed in prior research [68, 69, 70].
4.3 Verification Service
Stakeholders in Atlas seek to validate that artifacts have not been tampered with and that they were produced by expected pipelines running in high-integrity execution environments. To respond to verification requests, the Atlas verification service obtains golden values and transformation attestations from transparency logs relevant to an artifact in question.
First, the verification service validates the digital signatures on attestations and golden values to authenticate their producers. It then checks whether the artifact matches its golden value. If these checks pass, the service inspects the transformation attestations to confirm the ML system and pipeline operations ran as expected based on TEE attestations and golden values. Atlas validates artifact lineage by traversing the provenance chain, enabling efficient verification through batching related artifact types and maintaining a cache of verified transformations. This avoids repeated inspection of unchanged artifacts (R6), which particularly benefits iterative ML pipelines.
@@ 5 Implementation
Our proof-of-concept implementation integrates with PyTorch [38] and Kubeflow [37] through standard APIs for metadata tracking and execution monitoring within ML pipelines. This integration approach enabled us to avoid significant modifications to our case study pipelines (§6), while maintaining Atlas’ security and transparency enhancements. We leverage Intel TDX [10], a virtualization-layer TEE, to provide the hardware-based security primitives for ML systems and attestation clients in Atlas.
The attestation client is implemented as two components. First, a continuous ML system monitor integrates with PyTorch to collect metadata for a given pipeline. Second, the metadata sidecar (§5) running inside a dedicated Intel TDX TEE generates ML artifact and pipeline metadata in C2PA manifest format [16].
Due to current poor support for automated C2PA manifest generation for ML models, we implemented a Rust-based library and CLI [^4] Available at github.com/IntelLabs/atlas-cli that captures artifact measurements, Intel TDX attestations, and digital signatures in C2PA format. Supporting other software provenance formats [28, 29] is future work.
We extend Sigstore’s Rekor [32] to support Atlas C2PA-based model transformation attestations, validating signatures and measurements to ensure only properly signed artifacts are stored.
Out of space considerations, we provide additional details about the implementation in App. B.
Atlas Workflow Example. We illustrate Atlas’s end-to-end operations through an example with fine-tuning a BERT model for sentiment analysis:
  1. 1.
    Pipeline Environment Provisioning: MLaaS provider sets up Atlas attestation client and ML system monitor in Kubeflow.
  2. 2.
    Data Preparation: Data scientist prepares and uploads custom dataset, with Atlas metadata sidecar measuring and attesting the dataset using Intel TDX, submitting attestation to the transparency log.
  3. 3.
    Environment Validation: Atlas sidecar verifies training environment integrity, adding TDX-based attestation to the C2PA manifest.
  4. 4.
    Training Process Monitoring: Attestation client tracks (App. B.3):
    • Model weight changes via PyTorch hooks
    • Checkpoint creation and modifications
    • Hyperparameter updates
  5. 5.
    Model Verification: Third-party regulator verifies model provenance using Atlas verification service.
  6. 6.
    Deployment: Model vendor deploys verified model with provenance chain for user and application integrity validation.
@@
Metadata Sidecar. Because the Atlas ML system monitor runs alongside untrusted MLaaS provider code, the attestation client’s metadata sidecar leverages TEE remote attestation to detect tampering with the ML system monitor. That is, the sidecar interfaces with the ML system to obtain the Intel TDX-based compute environment attestations that capture TEE state and ML system component measurements, which are cryptographically anchored in hardware. We use the Confidential Containers (CoCo) framework [71] to implement the remote attestation procedure in the sidecar and ML system monitor.
Once the ML system monitor’s integrity has been validated, the sidecar generates and digitally signs C2PA manifests. These include the sidecar’s and ML system’s Intel TDX attestations, the received ML system metadata, the measurements for the pipeline’s input and output artifacts, and hashes for any linked transformation attestations. We describe a storage optimization in App. B.5.
Verification Service Implementation. For ease of implementation, the metadata sidecar also serves as a verification endpoint allowing pipeline components to validate artifact integrity against stored attestations. We optimize the performance of our staged verification system in three ways: 1) by processing changes incrementally and caching to avoid re-verifying unchanged components, 2) via batch processing of verification operations, and 3) parallel verification paths for independent component classes. App. B.6 provides additional details.
6 Evaluation
We validate our framework through a security analysis, preliminary performance testing and a case study with a BERT Meta [11, 12] fine-tuning pipeline.
6.1 Security Analysis
Atlas counters hub threats by verifying artifact integrity through cryptographic measurements and signatures, maintaining a provenance chain that identifies mismatched signatures or tampered dependencies before they propagate.
Atlas mitigates artifact producer threats through comprehensive provenance tracking, providing an immutable record of pipeline operations that detects undeclared dependencies and intentional omissions.
6.2 Preliminary Performance Analysis
We conduct our experiments on Intel® Xeon® Gold 5520+ processors with 256 GB of RAM running Ubuntu 24.04 beta. Employing Atlas with the BERT Meta case study’s CPU-only PyTorch-based fine-tuning pipeline, where the provenance chain covers $20$ artifacts (up to $120$ for more complex pipelines). Our measurements demonstrate near-linear scalability of verification time across different chain lengths and model sizes.
Preliminary tests show Atlas security mechanisms introduce minimal training overhead (under 8%), with each C2PA provenance manifest (8KB) containing artifact measurements, TEE attestations, and pipeline metadata. Verification processes scale linearly with model size, and our caching strategies reduce verification latency by up to 50%, achieving near-constant time for cached (see §5) component verification.
For large-scale operations, performance is maintained through concurrent verification operations, cache optimization, and selective invalidation for error handling. We plan to conduct more extensive benchmarking in future work.
6.3 Case Study
BERT was selected for its complex architecture and widespread production use. Our implementation covers the complete lifecycle from pre-trained model fine-tuning through deployment, with Intel TDX TEEs for attention computations and weight updates. Atlas secures instruction-based configuration using JSON records with query-positive-negative text triplets, tracking model adaptations and maintaining verifiable records of hyperparameters and training progressions.
The BERT Meta implementation demonstrates performance consistent with our analysis in §6.2.
6.4 Discussion & Limitations
Our implementation reveals limitations in current hardware security. While Atlas provides TEE-based protection for CPU operations [10], ML workloads rely on GPUs and accelerators without equivalent security features [72], creating security and trust boundaries between protected and unprotected environments [73]. Emerging solutions like confidential GPU computing show promise but have performance trade-offs [74, 75].
Additionally, ML lifecycle transparency creates competing requirements between verification and confidentiality of intellectual property and sensitive data [76], an important challenge that Atlas needs to address in a future version. Organizations must balance security requirements with operational efficiency, considering factors like verification frequency, attestation depth, and computational overhead.
7 Conclusion
The combination of hardware-backed security with runtime provenance tracking in Atlas provides a foundation for securing ML pipelines. Our case study shows Atlas’s ability to integrate into existing ML frameworks with reasonable performance. Interesting directions for future work include: 1. provenance tracking of ML accelerator-based computations, 2. end-to-end ML lifecycle confidentiality, and 3. algorithmic verification methods and model guardrails against attacks targeting model behavior.
Appendix A Background & Related Work
In addition to the works highlighted in §2, we describe further details and approaches about related work addressing ML lifecycle security and integrity.
A.1 Data Provenance & Authenticity
C2PA. The Coalition for Content Provenance and Authenticity (C2PA) specification [15, 16] was initially introduced by the Content Authenticity Initiative (CAI) [77] and Project Origin [78] as a response to the growing challenge of deepfakes [77] and digital content manipulation, gaining traction in digital photography and journalism workflows.
LakeFS. LakeFS [22] combines Git-like semantics with concepts from object stores such as S3 to provide a version control system for data, including ML datsets. Thus, LakeFS aims to capture data lineage by tracking changes to stored data over time, and allowing ML applications to reference specific versions of the stored data. This approach is meant to integrate with existing first-party data processing pipelines, but does not facilitate verification of data provenance by downstream consumers. Atlas’ metadata centered approach, on the other hand, enables first- and third-party ML dataset consumers to track changes and check their provenance, even when they may not have direct access to the data.
A.2 Supply Chain Integrity
BOM. Bills of Materials (BOM) have been employed to document the list of components of a hardware or software product for over three decades [79]. Software BOM (SBOM) have been the focus of many industry and academic efforts seeking to facilitate tracking software dependencies and other metadata [25], to improve their adoption, and to enhance SBOM integrity and privacy (e.g., [80, 81]).
Similarly, the AIBOM framework [26, 27] focuses on intended ML model supply chain management. Like SBOM, AIBOM provide a mechanism for tracking model software dependencies and maintaining model metadata. At the time of writing, we are not aware of any frameworks other than Atlas that utilize any sort of BOM data format to track ML model components.
Authenticated claims. A number of frameworks for capturing and verifying a variety of security claims and metadata about the supply chain have been proposed. in-toto [31] collects authenticated claims across supply chain steps, including SBOM and SLSA metadata. In particular, in-toto enables software development pipeline owners and downstream artifact consumers to specify end-to-end supply chain policies, and validate that only the expected parties carried out specific steps in the pipeline and artifacts underwent transformations in the expected order. Given recent and upcoming enhancements that further generalize the framework, in-toto may be a suitable option for specifying and verifying end-to-end ML model pipeline integrity policies in Atlas.
Sigstore [32] provides a transparency log-based infrastructure for issuing signing credentials and validating digital signatures on supply chain artifacts and metadata. Similarly, Supply Chain Integrity, Transparency and Trust (SCITT) [33] is an architecture for implementing distributed ledger based supply chain integrity mechanisms, providing global visibility and auditing for supply chain operations and claims. The SCITT architecture also includes confidential computing technologies that help ensure that only authorized parties submit claims to the transparency ledger.
Appendix B Implementation Details
B.1 Kubeflow Integration